1,3,2,6,4,9 what sequence is this

Answer by macston(5194)

.
Odd terms double, even terms add 3
.
ODD TERMS:
First term (1) x 2 = third term (2)
Third term (2) x 2 = fifth term (4)
Fifth term (4) x 2 = Seventh term would be 8
.
EVEN TERMS:
Second term(3) + 3 = fourth term (6)
Fourth term(6) + 3 = sixth term (9)
Sixth term (9) + 3 = eighth term would be 12